# Response of superfluid fermions at finite temperature

### v3

@@ II.3 The dynamical kernel
where the multi-indices $n=\{i_{n}f_{n}\}$ and $m=\{i_{m}f_{m}\}$ are introduced to simplify the notations, ${\bar{n}}=\{f_{n}i_{n}\}$, and the superfluid phonon vertices completed by the $\{{\cal U},{\cal V}\}$-amplitudes are specified in Appendix C. As in Ref. Litvinova and Zhang (2022), the additional index ”cc” marks the factorization with two correlation functions. The new features as compared to the zero-temperature case are the distributions $p_{i_{n}}$ and the exponential factors, in addition to the double summations over the initial states. Besides that, one finds the new terms with the $\{{\cal U},{\cal V}\}$ amplitudes, which appear with the $\Gamma^{(02)n}$ vertices, representing the diagonal part of the qPVC-associated ground state correlations. The ${\cal K}^{r[11]cc}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\omega_{k})$ of Eq. (115) can be used as the reference component for the entire dynamical kernel, so that the other components can be found analogously. For instance, in the ${\cal K}^{r[33]cc}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\omega_{k})$ matrix elements the amplitudes $\{{\cal X},{\cal Y}\}$ and $\{{\cal U},{\cal V}\}$ switch their roles as compared to the ${\cal K}^{r[11]cc}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\omega_{k})$ one, while ${\cal K}^{r[22]}$ and ${\cal K}^{r[44]}$ are related, respectively, to ${\cal K}^{r[11]}$ and ${\cal K}^{r[33]}$ by Hermitian conjugation. We note that the residues in Eq. (115) are nicely factorized so that block-non-diagonal matrix elements of the dynamical kernel can be found by recombining the amplitudes $\{{\cal X},{\cal Y},{\cal U},{\cal V}\}$ and vertices $\{\Gamma^{(11)},\Gamma^{(02)}\}$ according to their operator structure determined by Eq. (111). Furthermore, each type (a) component has a type (b) counterpart, where, in each residue, the indices $m$ and $n$ in the second pair of the transition amplitude and qPVC vertex interchange Litvinova and Zhang (2022).
@@ III Summary and outlook
We presented a response theory for superfluid fermionic systems at finite temperatures. The theory is formulated in the QFT language for fermions interacting via unspecified bare two-body forces. It starts with a definition of the two-quasiparticle superfluid thermal response function in the frequency domain, compatible with the structural part of the spectral strength distribution and with two-time two-fermion Matsubara propagator in a correlated medium averaged over the grand canonical ensemble in the imaginary time domain. The latter form enables a derivation of the most general EOM for this $2q$ propagator. As feedback, accounting for both forward and backward components of the response requires an additional thermal prefactor in the expression for the spectral strength distribution as compared to the zero-temperature case, which changes the low-energy limit of the strength distribution because of zero-energy singularity.
@@
The complete thermal superfluid formalism requires an extension of the regular zero-temperature 2$\times$2 component structure characterizing the fermionic response, to the 4$\times$4 arrays. This is dictated by the presence of transition amplitudes of the ${\cal U}_{\mu\mu^{\prime}}=\alpha^{\dagger}_{\mu}\alpha_{\mu^{\prime}}$ type, in addition to the commonly taken into account $\{{\cal X},{\cal Y}\}$ ones in QRPA and its extensions. Therefore, we included contributions from the ${\cal U}$-type amplitudes which vanish in QRPA at zero temperature, but acquire non-zero components in the norm matrix at finite temperature and thus become a non-negligible part of the EOM. This is a well-known feature of the FT-QRPA Sommermann (1983), which in this work is obtained in the FT-HFB limit of the static kernel and with the completely neglected dynamical kernel of the $2q$ EOM.
@@
The ${\cal U}$-type amplitudes contribute non-trivially to the dynamical kernel of the superfluid finite-temperature $2q$ EOM, i.e., in the FT-QRPA extension. Besides the formal extension of this kernel to the 4$\times$4 block form, these amplitudes further complicate the phonon vertex composition, appearing in the factorization approximations to the dynamical kernel which can be mapped to qPVC. Another new feature of the dynamical kernel is its explicit temperature dependence, in addition to that in the statistical thermal averages of all the expectation values, also present in the static kernel.
@@
This work thereby extends the applicability of the finite-temperature non-superfluid response theory formulated in Refs. Litvinova and Wibowo (2018); Wibowo and Litvinova (2019) and applied to various nuclear spectra characterizing both neutral and charge-exchange responses Litvinova and Wibowo (2018); Wibowo and Litvinova (2019); Litvinova et al. (2020); Litvinova and Robin (2021). The major new capability of the superfluid finite-temperature theory is bridging the superfluid phase at $T=0$ and the non-superfluid one at $T\geq T_{c}$, where $T_{c}$ is the critical temperature of the superfluid phase transition. The range $0\leq T\leq T_{c}$ is the most relevant temperature interval for nuclei undergoing rapid nucleosynthesis in stellar environments and for electron capture in core-collapse supernovae. Extending computation of these phenomena across the $0\leq T\leq T_{c}$ temperature regime is thereby one of the most interesting applications of the developed formalism to nuclear structure. Because of its generality, the theory can also be applied to a broad diapason of condensed matter phenomena.
@@ Acknowledgement
This work was partly supported by the US-NSF Grant PHY-2209376.
@@ Appendix B Generic commutators
$[V,A^{\dagger}_{\mu\mu^{\prime}}]=-[V,A_{\mu\mu^{\prime}}]^{\dagger},$ (129)
@@
Similarly to the A-block,
$[V,C^{\dagger}_{\mu\mu^{\prime}}]=-[V,C_{\mu\mu^{\prime}}]^{\dagger}.$ (137)
@@ Appendix C Phonon Vertices
-- $U^{T}\gamma^{n(-)T}U\Bigr{]}_{\mu\nu}$
@@
$\gamma^{n(+)}_{12}=\frac{1}{2}\sum\limits_{34}{\bar{v}}_{1234}\varkappa^{n(+)}_{34},\ \ \ \ \ \ \gamma_{12}^{n(-)T}=\frac{1}{2}\sum\limits_{34}{\bar{v}}^{\ast}_{1234}\varkappa^{n(-)\ast}_{34},$

### v4

@@ II.3 The dynamical kernel
where $\cal{AS}$ stands for the usual antisymmetrization $(\mu\leftrightarrow\mu^{\prime})$ and $(\nu\leftrightarrow\nu^{\prime})$, while $\cal{AS}^{\dagger}$ is defined in Appendix D. The multi-indices $n=\{i_{n}f_{n}\}$ and $m=\{i_{m}f_{m}\}$ are introduced to simplify the notations, ${\bar{n}}=\{f_{n}i_{n}\}$, and the superfluid phonon vertices completed by the $\{{\cal U},{\cal V}\}$-amplitudes are specified in Appendix C. As in Ref. Litvinova and Zhang (2022), the additional index ”cc” marks the factorization with two correlation functions. The new features, as compared to the zero-temperature case, are the distributions $p_{i_{n}}$ and the exponential factors, in addition to the double summations over the initial states. Besides that, one finds the new terms with the $\{{\cal U},{\cal V}\}$ amplitudes, which appear with the $\Gamma^{(02)n}$ vertices, representing the diagonal part of the qPVC-associated ground state correlations. The ${\tilde{\cal K}}^{r[11]cc}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\omega_{k})$ of Eq. (117) can be used as the reference component for the entire dynamical kernel, so that the other components can be found analogously. For instance, in the ${\tilde{\cal K}}^{r[33]cc}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\omega_{k})$ matrix elements the amplitudes $\{{\cal X},{\cal Y}\}$ and $\{{\cal U},{\cal V}\}$ switch their roles as compared to the ${\tilde{\cal K}}^{r[11]cc}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\omega_{k})$ one, while ${\tilde{\cal K}}^{r[22]}$ and ${\tilde{\cal K}}^{r[44]}$ are related, respectively, to ${\tilde{\cal K}}^{r[11]}$ and ${\tilde{\cal K}}^{r[33]}$ by Hermitian conjugation. We note that the residues in Eq. (117) are nicely factorized so that block-non-diagonal matrix elements of the dynamical kernel can be found by recombining the amplitudes $\{{\cal X},{\cal Y},{\cal U},{\cal V}\}$ and vertices $\{\Gamma^{(11)},\Gamma^{(02)}\}$ according to their operator structure determined by Eq. (111). Thus, generalizing Eqs. (117,119), one finds
${\tilde{\cal K}}^{r[ij]cc(a)}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\omega_{k})$ $=$ $\sum_{\gamma\delta nm}p_{i_{n}}p_{i_{m}}(1-e^{-(\omega_{n}+\omega_{m})/T})$
$\times$ $\Bigl[\frac{\langle[V,\hat{\Psi}^{[i]}_{\mu\mu^{\prime}}]\rangle^{nm}\langle[V,\hat{\Psi}^{[j]}_{\nu\nu^{\prime}}]\rangle^{nm\ast}}{\omega_{k}-\omega_{n}-\omega_{m}}$
- $\frac{\langle[V,\hat{\Psi}^{[i]}_{\mu\mu^{\prime}}]\rangle^{\bar{n}\bar{m}}\langle[V,\hat{\Psi}^{[j]}_{\nu\nu^{\prime}}]\rangle^{\bar{n}\bar{m}\ast}}{\omega_{k}+\omega_{n}+\omega_{m}}\Bigr],$
where the expectation values $\langle[V,\hat{\Psi}^{[i]}_{\mu\mu^{\prime}}]\rangle^{nm}$ are specified in Appendix D. Furthermore, each type (a) component has a type (b) counterpart, where, in each residue, the indices $m$ and $n$ in the second pair of the transition amplitude and qPVC vertex interchange Litvinova and Zhang (2022):
${\tilde{\cal K}}^{r[ij]cc(b)}_{\mu\mu^{\prime}\nu\nu^{\prime}}(\omega_{k})$ $=$ $\sum_{\gamma\delta nm}p_{i_{n}}p_{i_{m}}(1-e^{-(\omega_{n}+\omega_{m})/T})$ (121)
$\times$ $\Bigl[\frac{\langle[V,\hat{\Psi}^{[i]}_{\mu\mu^{\prime}}]\rangle^{nm}\langle[V,\hat{\Psi}^{[j]}_{\nu\nu^{\prime}}]\rangle^{mn\ast}}{\omega_{k}-\omega_{n}-\omega_{m}}$
- $\frac{\langle[V,\hat{\Psi}^{[i]}_{\mu\mu^{\prime}}]\rangle^{\bar{n}\bar{m}}\langle[V,\hat{\Psi}^{[j]}_{\nu\nu^{\prime}}]\rangle^{\bar{m}\bar{n}\ast}}{\omega_{k}+\omega_{n}+\omega_{m}}\Bigr].$
Thereby, Eqs. (117 - 121) represent the complete dynamical kernel in the ”cc”-factorization scheme. This factorization enables a closed EOM for the response function in terms of the two-quasiparticle amplitudes $\{{\cal X},{\cal Y}\}$ and $\{{\cal U},{\cal V}\}$, which can be solved by an iterative procedure, similarly to the non-superfluid case of Ref. Litvinova and Schuck (2019), confined by $\{{\cal X},{\cal Y}\}$. Strictly speaking, to have a fully closed matrix EOM with the bare interaction as the only input, the static kernel (157 - 162) should also be factorized.
Remarkably, here all the dynamical ground state correlations are included via the amplitudes $\{{\cal U},{\cal V}\}$ and off-diagonal components. The zero-temperature limit of Eqs. (117 - 121) can be easily taken, which means that the $T=0$ superfluid theory is also completed. Numerical implementations of the response theory with the dynamical kernels derived in this section will be explored for both $T=0$ and $T>0$ in future work.
@@ III Summary and outlook
We presented a response theory for superfluid fermionic systems at finite temperatures. The theory is formulated in the QFT language for fermions interacting via unspecified bare two-body forces. It starts with a definition of the two-quasiparticle superfluid thermal response function in the frequency domain, compatible with the structural part of the spectral strength distribution and with the two-time two-fermion Matsubara propagator in a correlated medium, averaged over the grand canonical ensemble in the imaginary time domain. The latter form enables a derivation of the most general EOM for this $2q$ propagator. As feedback, accounting for both forward and backward components of the response requires an additional thermal prefactor in the expression for the spectral strength distribution as compared to the zero-temperature case, which changes the low-energy limit of the strength distribution because of a zero-energy singularity.
@@
The complete thermal superfluid formalism requires an extension of the regular zero-temperature $\times2$ component structure characterizing the fermionic response to the $\times4$ arrays. This is dictated by the presence of transition amplitudes of the ${\cal U}_{\mu\mu^{\prime}}=\alpha^{\dagger}_{\mu}\alpha_{\mu^{\prime}}$ type, in addition to the commonly taken into account $\{{\cal X},{\cal Y}\}$ ones in QRPA and its extensions. Therefore, we included contributions from the $\cal U$}-type amplitudes, which vanish in QRPA at zero temperature, but acquire non-zero components in the norm matrix at finite temperature and thus become a non-negligible part of the EOM. This is a well-known feature of the FT-QRPA Sommermann (1983), which in this work is obtained in the FT-HFB limit of the static kernel and with the completely neglected dynamical kernel of the $2q$ EOM.
@@
The $\cal U$}-type amplitudes contribute non-trivially to the dynamical kernel of the superfluid finite-temperature $2q$ EOM, i.e., in the FT-QRPA extension. Besides the formal extension of this kernel to the $\times4$ block form, these amplitudes further complicate the phonon vertex composition, appearing in the factorization approximations to the dynamical kernel which can be mapped to qPVC. Another new feature of the dynamical kernel is its explicit temperature dependence, in addition to that in the statistical thermal averages of all the expectation values, also present in the static kernel.
@@
This work thereby extends the applicability of the finite-temperature non-superfluid response theory formulated in Refs. Litvinova and Wibowo (2018); Wibowo and Litvinova (2019) and applied to various nuclear spectra characterizing both neutral and charge-exchange responses Litvinova and Wibowo (2018); Wibowo and Litvinova (2019); Litvinova et al. (2020); Litvinova and Robin (2021). The major new capability of the superfluid finite-temperature theory is bridging the superfluid phase at $T=0$ and the non-superfluid one at $T\geq T_{c}$, where $T_{c}$ is the critical temperature of the superfluid phase transition. The range $0\leq T\leq T_{c}$ is the most relevant temperature interval for nuclei undergoing rapid nucleosynthesis in stellar environments and for electron capture in core-collapse supernovae. Extending computation of these phenomena across the $0\leq T\leq T_{c}$ temperature regime is thereby one of the most interesting applications of the developed formalism to nuclear structure. Because of its generality, the theory can also be applied across condensed matter phenomenology.
@@ Acknowledgement
This work was supported by the US-NSF Grants PHY-2209376 and PHY-2515056.
@@ Appendix B Generic commutators
$V,A^{\dagger}_{\mu\mu^{\prime}}]=-[V,A_{\mu\mu^{\prime}}]^{\dagger},$ (135)
@@
Similar to the A-block,
$V,C^{\dagger}_{\mu\mu^{\prime}}]=-[V,C_{\mu\mu^{\prime}}]^{\dagger}.$ (143)
@@ Appendix C Phonon Vertices
- $U^{T}\gamma^{n(-)T}U\Bigr]_{\mu\nu},$
$\Gamma^{(20)n}_{\mu\nu}=[U^{\dagger}g^{n}V^{\ast}$ $+$ $U^{\dagger}\gamma^{n(+)}U^{\ast}-V^{\dagger}g^{nT}U^{\ast}$ (147)
- $V^{\dagger}\gamma^{n(-)T}V^{\ast}\Bigr]_{\mu\nu},$
$\Gamma^{(20)n}_{\mu\nu}=-\Gamma^{(20)n}_{\nu\mu}=\Gamma^{(02)\bar{n}\ast}_{\mu\nu}=-\Gamma^{(02)\bar{n}\ast}_{\nu\mu},$ (148)
@@
$\gamma^{n(+)}_{12}$ $=$ $\frac{1}{2}\sum\limits_{34}{\bar{v}}_{1234}\varkappa^{n(+)}_{34}$,
$\gamma_{12}^{n(-)T}$ $=$ $\frac{1}{2}\sum\limits_{34}{\bar{v}}^{\ast}_{1234}\varkappa^{n(-)\ast}_{34}$, (151)
@@
Appendix D Expectation values and symmetry relations
The abbreviated forms (LABEL:Krijcca) and (121) of the (a) and (b) components of the factorized dynamical kernel, respectively, are defined via the following commutator expectation values:
$\langle[V,A_{\mu\mu^{\prime}}]\rangle^{nm}=\bigl[(\Gamma^{(11)n}_{\mu\gamma}{\cal X}^{m}_{\mu^{\prime}\gamma}+\Gamma^{(20)n}_{\mu\gamma}{\cal V}^{m}_{\mu^{\prime}\gamma}$
$-\frac{1}{2}\Gamma^{(20)n}_{\mu\mu^{\prime}}\delta_{i_{m}f_{m}})-(\mu\leftrightarrow\mu^{\prime})\bigr]=-\langle[V,A^{\dagger}_{\mu\mu^{\prime}}]\rangle^{{\bar{n}}{\bar{m}}\ast}$
and
$\langle[V,C_{\mu\mu^{\prime}}]\rangle^{nm}=\bigl[(\Gamma^{(11)n}_{\gamma\mu}{\cal U}^{m}_{\gamma\mu^{\prime}}+\Gamma^{(20)n}_{\mu^{\prime}\gamma}{\cal Y}^{m}_{\mu\gamma})$
$-(\mu\leftrightarrow\mu^{\prime})^{\dagger}\bigr]$
$=\bigl[(\Gamma^{(11){\bar{n}}\ast}_{\mu\gamma}{\cal U}^{{\bar{m}}\ast}_{\mu^{\prime}\gamma}-\Gamma^{(20){\bar{n}}\ast}_{\mu\gamma}{\cal Y}^{{\bar{m}}\ast}_{\mu^{\prime}\gamma})$
$-(\mu\leftrightarrow\mu^{\prime})^{\dagger}\bigr]=-\langle[V,C^{\dagger}_{\mu\mu^{\prime}}]\rangle^{{\bar{n}}{\bar{m}}\ast},$ (153)
where
$(\mu\leftrightarrow\mu^{\prime})^{\dagger}=\left(\begin{array}[]{c}\mu\leftrightarrow\mu^{\prime}\\ n\leftrightarrow{\bar{n}}\\ m\leftrightarrow{\bar{m}}\end{array}\right)={\cal AS}^{\dagger}.$ (154)
The following symmetry relations between the transition density components were employed:
${\cal X}_{\mu\mu^{\prime}}^{n}$ $\equiv$ $\langle i|\alpha_{\mu^{\prime}}\alpha_{\mu}|f\rangle=\langle f|\alpha_{\mu}^{\dagger}\alpha_{\mu^{\prime}}^{\dagger}|i\rangle^{\ast}\equiv{\cal Y}_{\mu\mu^{\prime}}^{{\bar{n}}\ast}$ (155)
${\cal U}_{\mu\mu^{\prime}}^{n}$ $\equiv$ $\langle i|\alpha_{\mu}^{\dagger}\alpha_{\mu^{\prime}}|f\rangle=\langle f|\alpha_{\mu^{\prime}}^{\dagger}\alpha_{\mu}|i\rangle^{\ast}\equiv{\cal V}_{\mu\mu^{\prime}}^{{\bar{n}}\ast}.$ (156)
Appendix E Double commutators of the static kernel
The diagonal matrix elements of the static kernel read:
$T^{0[11]}_{\mu\mu^{\prime}\nu\nu^{\prime}}$ $=$ $-\langle[[V,\Psi^{[1]}_{\mu\mu^{\prime}}],\Psi^{[1]\dagger}_{\nu\nu^{\prime}}]\rangle=-\langle[[V,A_{\mu\mu^{\prime}}],A^{\dagger}_{\nu\nu^{\prime}}]\rangle$ (157)
$=$ $-\langle[[V,\Psi^{[2]}_{\mu\mu^{\prime}}],\Psi^{[2]\dagger}_{\nu\nu^{\prime}}]\rangle^{\ast}=-\langle[[V,A^{\dagger}_{\mu\mu^{\prime}}],A_{\nu\nu^{\prime}}]\rangle^{\ast}$
$=$ $T^{0[22]*}_{\mu\mu^{\prime}\nu\nu^{\prime}}$,
$T^{0[11]}_{\mu\mu^{\prime}\nu\nu^{\prime}}$ $=$ $-\langle[[V,A_{\mu\mu^{\prime}}],A^{\dagger}_{\nu\nu^{\prime}}]\rangle$ (158)
$=$ [{ $\Biggl[\Bigl\{\,2\,H^{40}_{\rho\mu\gamma\gamma^{\prime}}\bigl({\cal D}^{[21]}_{\nu\rho\gamma\gamma^{\prime}}+{\cal D}^{[21]}_{\gamma\gamma^{\prime}\nu\rho}\bigr)\delta_{\nu^{\prime}\mu^{\prime}}$
$+$ $H^{31}_{\rho\rho^{\prime}\mu\gamma}\Bigl(\bigl({\cal D}^{[24]}_{\rho\rho^{\prime}\nu^{\prime}\gamma}\,\delta_{\mu^{\prime}\nu}-(\mu^{\prime}\leftrightarrow\gamma)\bigr)$
- $2\bigl({\cal D}^{[24]}_{\nu\rho\rho^{\prime}\gamma}\delta_{\nu^{\prime}\mu^{\prime}}+{\cal D}^{[41]}_{\mu^{\prime}\rho\nu\rho^{\prime}}\,\delta_{\nu^{\prime}\gamma}\bigr)\Bigr)$
$+$ $H^{31*}_{\rho\rho^{\prime}\gamma\mu}\Bigl(\bigl({\cal D}^{[32]}_{\nu^{\prime}\gamma\rho\rho^{\prime}}\,\delta_{\mu^{\prime}\nu}-(\mu^{\prime}\leftrightarrow\gamma)\bigr)$
$+$ $\bigl({\cal D}^{[14]}_{\mu^{\prime}\gamma\nu^{\prime}\rho^{\prime}}\,\delta_{\rho\nu}-(\rho\leftrightarrow\rho^{\prime})\bigr)\Bigr)$
$+$ $H^{22}_{\gamma\mu\rho\rho^{\prime}}\Bigl({\cal D}^{[34]}_{\gamma\mu^{\prime}\nu^{\prime}\rho^{\prime}}\,\delta_{\rho\nu}-(\rho\leftrightarrow\rho^{\prime})\Bigr)$
$+$ $H^{22}_{\gamma\mu\rho\rho^{\prime}}\,{\cal D}^{[22]}_{\nu\gamma\rho\rho^{\prime}}\,\delta_{\nu^{\prime}\mu^{\prime}}\Bigr\}-\{\mu\leftrightarrow\mu^{\prime}\}\Biggr]$
- $[\nu\leftrightarrow\nu^{\prime}],$
$T^{0[33]}_{\mu\mu^{\prime}\nu\nu^{\prime}}$ $=$ $-\langle[[V,\Psi^{[3]}_{\mu\mu^{\prime}}],\Psi^{[3]\dagger}_{\nu\nu^{\prime}}]\rangle=-\langle[[V,C_{\mu\mu^{\prime}}],C^{\dagger}_{\nu\nu^{\prime}}]\rangle$ (159)
$=$ $-\langle[[V,\Psi^{[4]}_{\mu\mu^{\prime}}],\Psi^{[4]\dagger}_{\nu\nu^{\prime}}]\rangle^{\ast}=-\langle[[V,C^{\dagger}_{\mu\mu^{\prime}}],C_{\nu\nu^{\prime}}]\rangle^{\ast}$
$=$ $T^{0[44]*}_{\mu\mu^{\prime}\nu\nu^{\prime}}$,
$T^{0[33]}_{\mu\mu^{\prime}\nu\nu^{\prime}}$ $=$ $-\langle[[V,C_{\mu\mu^{\prime}}],C^{\dagger}_{\nu\nu^{\prime}}]\rangle$ (162)
$=$ $\Bigl[4H^{40}_{\rho\rho^{\prime}\gamma\mu^{\prime}}\Big({\cal D}^{[21]}_{\rho\nu^{\prime}\gamma\mu}\,\delta_{\rho^{\prime}\nu}-{\cal D}^{[21]}_{\rho^{\prime}\nu^{\prime}\gamma\mu}\,\delta_{\rho\nu}$
$\;+{\cal D}^{[21]}_{\rho\rho^{\prime}\gamma\nu^{\prime}}\,\delta_{\mu\nu}-{\cal D}^{[21]}_{\rho\rho^{\prime}\mu\nu^{\prime}}\,\delta_{\gamma\nu}\Big)$
$\;-H^{31}_{\rho\rho^{\prime}\gamma\mu}\Big({\cal D}^{[24]}_{\rho^{\prime}\nu^{\prime}\gamma\mu^{\prime}}\,\delta_{\rho\nu}-{\cal D}^{[24]}_{\rho\nu^{\prime}\gamma\mu^{\prime}}\,\delta_{\rho^{\prime}\nu}$
$\;+{\cal D}^{[24]}_{\rho\rho^{\prime}\gamma\nu}\,\delta_{\mu^{\prime}\nu^{\prime}}-{\cal D}^{[23]}_{\rho\rho^{\prime}\mu^{\prime}\nu^{\prime}}\,\delta_{\gamma\nu}\Big)$
$\;+H^{31}_{\rho\rho^{\prime}\mu^{\prime}\gamma}\Big({\cal D}^{[23]}_{\rho^{\prime}\nu^{\prime}\gamma\mu}\,\delta_{\rho\nu}-{\cal D}^{[23]}_{\rho\nu^{\prime}\gamma\mu}\,\delta_{\rho^{\prime}\nu}$
$\;+{\cal D}^{[24]}_{\rho\rho^{\prime}\mu\nu}\,\delta_{\gamma\nu^{\prime}}-{\cal D}^{[23]}_{\rho\rho^{\prime}\gamma\nu^{\prime}}\,\delta_{\mu\nu}\Big)$
$\;-2H^{31}_{\gamma\mu^{\prime}\rho\rho^{\prime}}\Big({\cal D}^{[24]}_{\mu\nu^{\prime}\rho\rho^{\prime}}\,\delta_{\gamma\nu}-{\cal D}^{[24]}_{\gamma\nu^{\prime}\rho\rho^{\prime}}\,\delta_{\mu\nu}$
$\;+{\cal D}^{[24]}_{\gamma\mu\rho\nu}\,\delta_{\rho^{\prime}\nu^{\prime}}-{\cal D}^{[23]}_{\gamma\mu\rho^{\prime}\nu^{\prime}}\,\delta_{\rho\nu}\Big)$
$\;-\frac{1}{2}\,H^{22}_{\gamma\mu^{\prime}\rho\rho^{\prime}}\Big({\cal D}^{[22]}_{\mu\nu^{\prime}\rho\rho^{\prime}}\,\delta_{\gamma\nu}-{\cal D}^{[22]}_{\gamma\nu^{\prime}\rho\rho^{\prime}}\,\delta_{\mu\nu}$
$\;+{\cal D}^{[22]}_{\gamma\mu\rho\nu}\,\delta_{\rho^{\prime}\nu^{\prime}}-{\cal D}^{[22]}_{\gamma\mu\rho^{\prime}\nu}\,\delta_{\rho\nu^{\prime}}\Big)\Bigr]$
$\;-\Bigl[\begin{array}[]{c}\mu\leftrightarrow\mu^{\prime}\\ \nu\leftrightarrow\nu^{\prime}\end{array}\Bigr]^{\ast},$
where summation is implied over the repeated indices, and the generalized superfluid two-body density is used:
${\cal D}_{\mu\mu^{\prime}\nu\nu^{\prime}}$ $=$ $\left\langle\left(\begin{array}[]{cccc}A_{\mu\mu^{\prime}}A^{\dagger}_{\nu\nu^{\prime}}&A_{\mu\mu^{\prime}}A_{\nu\nu^{\prime}}&A_{\mu\mu^{\prime}}C^{\dagger}_{\nu\nu^{\prime}}&A_{\mu\mu^{\prime}}C_{\nu\nu^{\prime}}\\[4.0pt] A^{\dagger}_{\mu\mu^{\prime}}A^{\dagger}_{\nu\nu^{\prime}}&A^{\dagger}_{\mu\mu^{\prime}}A_{\nu\nu^{\prime}}&A^{\dagger}_{\mu\mu^{\prime}}C^{\dagger}_{\nu\nu^{\prime}}&A^{\dagger}_{\mu\mu^{\prime}}C_{\nu\nu^{\prime}}\\[4.0pt] C_{\mu\mu^{\prime}}A^{\dagger}_{\nu\nu^{\prime}}&C_{\mu\mu^{\prime}}A_{\nu\nu^{\prime}}&C_{\mu\mu^{\prime}}C^{\dagger}_{\nu\nu^{\prime}}&C_{\mu\mu^{\prime}}C_{\nu\nu^{\prime}}\\[4.0pt] C^{\dagger}_{\mu\mu^{\prime}}A^{\dagger}_{\nu\nu^{\prime}}&C^{\dagger}_{\mu\mu^{\prime}}A_{\nu\nu^{\prime}}&C^{\dagger}_{\mu\mu^{\prime}}C^{\dagger}_{\nu\nu^{\prime}}&C^{\dagger}_{\mu\mu^{\prime}}C_{\nu\nu^{\prime}}\end{array}\right)\right\rangle$ (167)
$=$ $\left\langle\left(\begin{array}[]{c}A_{\mu\mu^{\prime}}\\[2.0pt] A^{\dagger}_{\mu\mu^{\prime}}\\[2.0pt] C_{\mu\mu^{\prime}}\\[2.0pt] C^{\dagger}_{\mu\mu^{\prime}}\end{array}\right)\!\!\left(\begin{array}[]{cccc}A^{\dagger}_{\nu\nu^{\prime}}&A_{\nu\nu^{\prime}}&C^{\dagger}_{\nu\nu^{\prime}}&C_{\nu\nu^{\prime}}\end{array}\right)\right\rangle.$ (173)